I am having trouble with this equation problem : 2z+4=5z-8
I was taught to isolate the variables,
i.e. 2z+4-4=5z-8-4
2z=5z-4
But sometimes i don't know which part of the question I should isolate first.
Can you explain to me the formula/rule?

Answer by checkley71(8403)

2Z+4-4=5Z-8-4 OR 2Z-5Z=-8-4-4+4 OR -3Z=-12 OR 3Z=12 OR Z=12/3 OR Z=4